# Burnside type results for fusion categories

### v10

@@ 2.4. Grouplike elements in hypergroups
In any symmetric RN-hypergroup ${(H,\mathcal{B})}$, the set $\{{\frac{x_{i}}{d_{i}}}\}$ of normalized grouplike-elements form a finite group. Moreover, the inverse of ${\frac{x_{i}}{d_{i}}}$ is given by $\frac{x_{i^{*}}}{d_{i^{*}}}$ for all $x_{i}\in{G{(H,\mathcal{B})}}$.
@@
If $x_{i}x_{i^{*}}=\frac{1}{h_{i}}x_{1}$ and $x_{j}x_{j^{*}}=\frac{1}{h_{j}}x_{1}$ it follows that $(x_{i}x_{j})(x_{i}x_{j})^{*}=\frac{1}{h_{i}h_{j}}x_{1}.$ Thus $(\sum_{k=1}^{m}N^{k}_{ij}x_{k})(\sum_{k=1}^{m}N^{k}_{ij}x_{k^{*}})=\frac{1}{h_{i}h_{j}}x_{1}$ and this implies that $x_{i}x_{j}=\frac{d_{j}d_{i}}{d_{k}}x_{k}$ for a certain $k\in\mathcal{J}$. Moreover, since $h_{i}=\frac{1}{d_{i}^{2}}$ and $h_{j}=\frac{1}{d_{j}^{2}}$ it follows that
$h_{k}^{-1}=m(x_{1},x_{k}x_{k}^{*})=\frac{d_{k}^{2}}{d_{i}^{2}d_{j}^{2}}m(x_{1},x_{i}x_{j}x_{i^{*}}x_{j^{*}})=\frac{d_{k}^{2}}{d_{i}^{2}d_{j}^{2}}\frac{1}{h_{i}h_{j}}=d_{k}^{2}$
which shows that $x_{k}$ is also a grouplike element ($m$ is defined in §5.1). ∎
@@
By duality one obtains the following:
@@
If ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup and the dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN then ${G{\widehat{(H,\mathcal{B})}}}$ is a group.
@@
Let ${(H,\mathcal{B})}$ be a symmetric RN-hypergroup. If $x_{i}\in{G{(H,\mathcal{B})}}$ is a grouplike element then ${\frac{x_{i}}{d_{i}}}\frac{x_{j}}{d_{j}},\frac{x_{j}}{d_{j}}{\frac{x_{i}}{d_{i}}}\in\overline{\mathcal{B}}$, for any $x_{j}\in\mathcal{B}$.
@@
where $\mathcal{A}$ is a set such that $N^{k}_{ij}>0$. Multiplying the above equality by $x_{i^{*}}$, one obtains
@@
since $h_{i}x_{i^{*}}x_{i}=x_{1}$ as $x_{i}$ is a grouplike element. Since ${(H,\mathcal{B},\mu_{1})}$ is a RN-hypergroup, it follows that $x_{i^{*}}x_{k}=\alpha_{ik}x_{j}$, for some scalar $\alpha_{ik}\in\mathbb{C}$ and all $k\in\mathcal{A}$. Thus $\alpha_{ik}=\frac{d_{i}d_{k}}{d_{j}}$. Moreover, multiplying the last equality by $x_{i}$, one obtains that $\frac{1}{h_{i}}x_{k}=x_{i^{*}}x_{j}$, which shows that $\mathcal{A}$ has a single element. ∎
@@
Suppose that ${(H,\mathcal{B})}$ is an abelian RN hypergroup and $\mu_{j}\in{\widehat{(H,\mathcal{B})}}$ a character such that $\mu_{j}(x_{m})=d_{m}\omega_{m}$ and $\mu_{j}(x_{n})=d_{n}\omega_{n}$ for some scalars $|\omega_{m}|=|\omega_{n}|=1$. Then $\mu_{j}(x_{p})=\omega_{m}\omega_{n}d_{p}$ for all constituents $x_{p}$ of $x_{m}x_{n}$.
@@
Straightforward from the absolute value triangle inequality. ∎
@@
If ${(H,\mathcal{B})}$ is an abelian RN hypergroup then ${{G{\widehat{(H,\mathcal{B})}}}}$ is a group.
@@
Suppose that $\mu_{j},\mu_{k}\in{{G{(H,\mathcal{B})}}}$. By Lemma 2.21 it follows that $\mu_{j}(x_{m})=\omega_{m}\mathrm{FPdim}(x_{m})$ for some roots of unity $\omega_{m}$. Similarly $\mu_{k}(x_{m})=\eta_{m}\mathrm{FPdim}(x_{m})$ for some other root of unity $\eta_{m}$. Moreover, by lemma 2.28, since ${(H,\mathcal{B})}$ is RN it follows that all the constituents $x_{m}x_{n}$ when evaluated at $\mu_{j}$ have the scalar $\omega_{m}\omega_{n}$. On the other hand, evaluated at $\mu_{k}$ these constituents give the scalar $\eta_{m}\eta_{n}$. Thus
$(\mu_{j}\star\mu_{k})(x_{m}x_{n})$ $=$ $\sum_{p}N^{p}_{mn}\mathrm{FPdim}(x_{p})(\mu_{j}\star\mu_{k})({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})$
$=$ $\sum_{p}N^{p}_{mn}\mathrm{FPdim}(x_{p})\mu_{j}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})\mu_{k}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})$
$=$ $\omega_{m}\omega_{n}\eta_{m}\eta_{n}\mathrm{FPdim}(x_{m}x_{n})$
On the other hand, note that $\mu_{j}(x_{m})=\mathrm{FPdim}(x_{m})\mu_{j}({\frac{x_{m}}{\mathrm{FPdim}(x_{m})}})=\omega_{m}\mathrm{FPdim}(x_{m})$ and similarly $\mu_{k}(x_{m})=\mathrm{FPdim}(x_{m})\mu_{k}({\frac{x_{m}}{\mathrm{FPdim}(x_{m})}})=\eta_{m}\mathrm{FPdim}(x_{m}).$ Therefore
$(\mu_{j}\star\mu_{k})(x_{m})(\mu_{j}\star\mu_{k})(x_{n})=\omega_{m}\eta_{m}\mathrm{FPdim}(x_{m})\omega_{n}\eta_{n}\mathrm{FPdim}(x_{n})$
This shows that $(\mu_{j}\star\mu_{k})(x_{m}x_{n})=(\mu_{j}\star\mu_{k})(x_{m})(\mu_{j}\star\mu_{k})(x_{n})$ which shows that $(\mu_{j}\star\mu_{k})\in{\widehat{\mathcal{B}}}$. Moreover, $(\mu_{j}\star\mu_{k})({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})=\mu_{j}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})\mu_{k}({\frac{x_{p}}{\mathrm{FPdim}(x_{p})}})=\omega_{p}\eta_{p}$ and Lemma 2.21 implies that $\mu_{j}\star\mu_{k}\in G{\widehat{(H,\mathcal{B})}}$. ∎
@@
Let ${(H,\mathcal{B})}$ be an abelian RN hypergroup and $\mu\in{{G{\widehat{(H,\mathcal{B})}}}}$. It follows that $\mu\star\mu_{k}\in\widehat{B}$ for any $\mu_{k}\in\widehat{B}$.
@@
Since $\mu\in{{G{\widehat{(H,\mathcal{B})}}}}$ and ${(H,\mathcal{B})}$ is RN, by Lemma 2.21 one has $\mu(x_{i})=\omega_{i}d_{i}$ for some root of unity $\omega_{i}$ and for all $i\in\mathcal{I}$. Here $d_{i}=\mathrm{FPdim}(x_{i})$. Then as above it follows that $\mu(x_{k})=\omega_{i}\omega_{j}d_{k}$ for any constituent $x_{k}$ of $x_{i}x_{j}$.
One can compute directly that $\mu\star\mu_{k}$ and $\mu_{k}\star\mu$ are morphisms of algebras. Indeed, note that $(\mu\star\mu_{k})({\frac{x_{i}}{d_{i}}})=\omega_{i}\mu_{k}({\frac{x_{i}}{d_{i}}})$ and therefore
@@
In the dual situation, Lemma 2.29 and Proposition 2.30 can be written as follows:
@@
If ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN then $G{(H,\mathcal{B})}$ is a group. Moreover, left multiplication by $L_{{\frac{x_{i}}{d_{i}}}}$ for any $x_{i}\in{G{(H,\mathcal{B})}}$ permutes the normalized standard basis $\overline{\mathcal{B}}$.
Note that Lemmata 2.25, 2.26, 2.27, 2.29 and Propositions 2.30, 2.31 imply the following:
@@
    ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
Then the following statements hold:
@@
    The operator given by left multiplication $L_{\frac{x_{i}}{d_{i}}}$ with $x_{i}\in G{(H,\mathcal{B})}$ permutes the normalized standard basis $\overline{\mathcal{B}}$.
@@
    The operator given by left multiplication $L_{\mu_{j}}$ with $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ permutes the standard basis $^{{\widehat{\mathcal{B}}}}$.
2.5. Some results on $P$ and ${\widehat{P}}$.
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup, and let $x_{i}\in G{(H,\mathcal{B})}$ be a grouplike element. Then
(2.24) $<\prod_{j=1}^{m}\mu_{j},\;{\frac{x_{i}}{d_{i}}}>=\pm 1.$
@@
By the definition of the multiplication in the dual,
$<\prod_{j=1}^{m}\mu_{j},\;{\frac{x_{i}}{d_{i}}}>=\prod_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}}).$
We group together the factors of $\mu_{j}$ and $\mu_{j\#}$ in the above product. Note that $\mu_{j}({\frac{x_{i}}{d_{i}}})\mu_{j^{\#}}({\frac{x_{i}}{d_{i}}})=\mid\mu_{j}({\frac{x_{i}}{d_{i}}})\mid^{2}=1$ by Lemma 2.20. On the other hand, if $\mu_{j}=\mu_{j^{\#}}$, then $\mu_{j}({\frac{x_{i}}{d_{i}}})\in\mathbb{R}$ and therefore $\mu_{j}({\frac{x_{i}}{d_{i}}})=\pm 1$ by the same Lemma 2.20. ∎
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup and let $x_{i}\in{G{(H,\mathcal{B})}}$ be a grouplike element. We denote by ${\mathrm{sgn}}(x_{i})$ the above value of the product $\prod_{j=1}^{m}\mu_{j}$ evaluated at ${\frac{x_{i}}{d_{i}}}$.
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup, and let $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ be a grouplike element. Then
(2.25) $<\mu_{j},\;{\prod_{i=1}^{m}}{\frac{x_{i}}{d_{i}}}>=\pm 1.$
By duality we can also define:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup and let $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ be a grouplike element. We denote by ${\mathrm{sgn}}(\mu_{j})$ the above value of the product ${\prod_{i=1}^{m}}{\frac{x_{i}}{d_{i}}}$ evaluated by $\mu_{j}$.
Remark 2.37.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalized hypergroup and let $x_{i}\in{G{(H,\mathcal{B})}}$ be a grouplike element. Then $L_{{\frac{x_{i}}{d_{i}}}}$, the left multiplication by ${\frac{x_{i}}{d_{i}}}$ on $H$ induces a permutation on $\overline{\mathcal{B}}$. Because ${\frac{x_{i}}{d_{i}}}=\sum_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}})F_{j}$, the determinant $\det(L_{\frac{x_{i}}{d_{i}}})$ of the matrix of the operator $L_{{\frac{x_{i}}{d_{i}}}}$ is $\prod_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}})$. On the other hand, $L_{{\frac{x_{i}}{d_{i}}}}$ permutes the normalized basis ¯:=$\overline{\mathcal{B}}:=\{{\frac{x_{i}}{d_{i}}}\}_{i\in\mathcal{I}}$, and therefore, ${\mathrm{sgn}}(x_{i})$ is the signature of the corresponding permutation.
Remark 2.38.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup and $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$. It follows that ${\mathrm{sgn}}(\mu_{j})=\det(L_{\mu_{j}})=\pm 1$ since $\mu_{j}$ determines a permutation of $^{\widehat{\mathcal{B}}}$.
@@
A symmetric hypergroup ${(H,\mathcal{B})}$ is called perfect if it has no nontrivial grouplike elements.

### v11

@@ 2.5. Grouplike elements in hypergroups
In any RN-hypergroup ${(H,\mathcal{B})}$, the set $\{{\frac{x_{i}}{d_{i}}}\}$ of normalized grouplike-elements forms a finite group. Moreover, the inverse of ${\frac{x_{i}}{d_{i}}}$ is given by $\frac{x_{i^{*}}}{d_{i^{*}}}$ for all $x_{i}\in{G{(H,\mathcal{B})}}$.
@@
If $x_{i}x_{i^{*}}=\frac{1}{h_{i}}x_{1}$ and $x_{j}x_{j^{*}}=\frac{1}{h_{j}}x_{1}$ then $(x_{i}x_{j})(x_{i}x_{j})^{*}=\frac{1}{h_{i}h_{j}}x_{1}.$ Thus $(\sum_{k=1}^{m}N^{k}_{ij}x_{k})(\sum_{k=1}^{m}N^{k}_{ij}x_{k^{*}})=\frac{1}{h_{i}h_{j}}x_{1},$ so that there is a unique $k\in\mathcal{I}$ such that $N_{i,j}^{k}$ is nonzero (if it were not the case, then by the axioms of hypergroups coupled with the RN assumption, the left-hand side of the aforementioned equation would decompose into several components). Consequently, we have $(N^{k}_{ij}x_{k})(N^{k}_{ij}x_{k^{*}})=\frac{1}{h_{i}h_{j}}x_{1}$, and furthermore, $x_{i}x_{j}=N^{k}_{ij}x_{k}$. However, applying $\mathrm{FPdim}$ yields $N^{k}_{ij}=(d_{i}d_{j})/d_{k}$, so $\frac{x_{i}}{d_{i}}\frac{x_{j}}{d_{j}}=\frac{x_{k}}{d_{k}}$, and by the definition of a grouplike element, we have $h_{i}^{-1}=d_{i}^{2}$. Synthesizing all these equalities, we deduce that $((d_{i}d_{j})/d_{k})^{2}x_{k}x_{k^{*}}=(d_{i}d_{j})^{2}x_{1}$, leading to the conclusion that $x_{k}x_{k^{*}}=d_{k}^{2}x_{1}=h_{k}^{-1}x_{1}$. There remain to prove that if $x_{i}\in{G{(H,\mathcal{B})}}$ then so is $x_{i^{*}}$. By Lemma 2.24, $x_{i^{*}}$ is grouplike if and only if $h_{i^{*}}d_{i^{*}}^{2}=1$. But $h_{i^{*}}=h_{i}$ by Theorem 2.16, and $d_{i^{*}}=d_{i}$ by Corollary 2.18, so $h_{i^{*}}d_{i^{*}}^{2}=h_{i}d_{i}^{2}=1$, because $x_{i}$ is grouplike. Finally, by above, there is $k$ such that $\frac{x_{i}}{d_{i}}\frac{x_{i^{*}}}{d_{i^{*}}}=\frac{x_{k}}{d_{k}}$, and by the hypergroup axioms, $k$ must be 111. ∎
@@
By duality we obtain the following:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that its dual $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then $^{G{\widehat{(H,\mathcal{B})}}}$ is a group.
@@
Let ${(H,\mathcal{B})}$ be a RN-hypergroup. Let $x_{i}\in{G{(H,\mathcal{B})}}$, then ${\frac{x_{i}}{d_{i}}}\frac{x_{j}}{d_{j}},\frac{x_{j}}{d_{j}}{\frac{x_{i}}{d_{i}}}\in\overline{\mathcal{B}}$, for any $x_{j}\in\mathcal{B}$.
@@
where $\mathcal{A}$ is a set such that $N^{k}_{ij}>0$. Multiplying the above equality by $x_{i^{*}}$, we obtain
@@
since $h_{i^{*}}x_{i^{*}}x_{i}=x_{1}$ as $x_{i^{*}}$ is also grouplike by Lemma 2.32, and $h_{i^{*}}=h_{i}$ by Theorem 2.16. By RN assumption, $x_{i^{*}}x_{k}=\alpha_{ik}x_{j}$, for some scalar $\alpha_{ik}>0$ and all $k\in\mathcal{A}$. Thus $\alpha_{ik}=\frac{d_{i}d_{k}}{d_{j}}$. Multiplying the second last equality by $x_{i}$, we get that $h_{i}^{-1}x_{k}=\alpha_{ik}x_{i}x_{j}$, which reformulates as ${\frac{x_{i}}{d_{i}}}\frac{x_{j}}{d_{j}}=\frac{x_{k}}{d_{k}}\in\overline{\mathcal{B}}$. Idem for $\frac{x_{j}}{d_{j}}{\frac{x_{i}}{d_{i}}}$ using right multiplications. ∎
@@
Suppose that ${(H,\mathcal{B})}$ is an abelian RN hypergroup and $\mu_{j}\in\widehat{\mathcal{B}}$ a character such that $\mu_{j}(x_{m})=d_{m}\omega_{m}$ and $\mu_{j}(x_{n})=d_{n}\omega_{n}$ for some scalars $|\omega_{m}|=|\omega_{n}|=1$. Then $\mu_{j}(x_{p})=\omega_{m}\omega_{n}d_{p}$ for all constituents $x_{p}$ of $x_{m}x_{n}$.
@@
Apply the absolute value triangle inequality and Lemma 2.30. ∎
@@
If ${(H,\mathcal{B})}$ is an abelian RN hypergroup then $^{{G{\widehat{(H,\mathcal{B})}}}}$ is a group.
@@
Suppose that $\mu_{j},\mu_{k}\in{{G{\widehat{(H,\mathcal{B})}}}}$. By Lemma 2.28, $\mu_{j}(x_{m})=\omega_{m}d_{m}$ and $\mu_{k}(x_{m})=\eta_{m}d_{m}$, for some roots of unity $\omega_{m}$ and $\eta_{m}$. Now,
$(\mu_{j}\star\mu_{k})(x_{m}x_{n})=\sum_{p}N^{p}_{m,n}d_{p}(\mu_{j}\star\mu_{k})({\frac{x_{p}}{d_{p}}}),$
but by Equation (2.7),
$(\mu_{j}\star\mu_{k})({\frac{x_{p}}{d_{p}}})=\mu_{j}({\frac{x_{p}}{d_{p}}})\mu_{k}({\frac{x_{p}}{d_{p}}}),$
and by Lemma 2.35, $\mu_{j}(x_{p})=\omega_{m}\omega_{n}d_{p}$ and $\mu_{k}(x_{p})=\eta_{m}\eta_{n}d_{p}$, when $N^{p}_{m,n}$ is nonzero. Thus
$(\mu_{j}\star\mu_{k})(x_{m}x_{n})=\sum_{p}N^{p}_{m,n}d_{p}\omega_{m}\omega_{n}\eta_{m}\eta_{n}=\omega_{m}\omega_{n}\eta_{m}\eta_{n}d_{m}d_{n}$
On the other hand,
$(\mu_{j}\star\mu_{k})(x_{m})=d_{m}(\mu_{j}\star\mu_{k})({\frac{x_{m}}{d_{m}}})=d_{m}\mu_{j}({\frac{x_{m}}{d_{m}}})\mu_{k}({\frac{x_{m}}{d_{m}}})=d_{m}\omega_{m}\eta_{m}.$
Idem, $(\mu_{j}\star\mu_{k})(x_{n})=d_{n}\omega_{n}\eta_{n}$. So
$(\mu_{j}\star\mu_{k})(x_{m})(\mu_{j}\star\mu_{k})(x_{n})=d_{m}\omega_{m}\eta_{m}d_{n}\omega_{n}\eta_{n}=(\mu_{j}\star\mu_{k})(x_{m}x_{n}),$
by above, which shows that $(\mu_{j}\star\mu_{k})\in{\widehat{\mathcal{B}}}$.
Moreover, $(\mu_{j}\star\mu_{k})({\frac{x_{p}}{d_{p}}})=\mu_{j}({\frac{x_{p}}{d_{p}}})\mu_{k}({\frac{x_{p}}{d_{p}}})=\omega_{p}\eta_{p}$, which by Lemma 2.28, implies that $\mu_{j}\star\mu_{k}\in G{\widehat{(H,\mathcal{B})}}$. ∎
@@
Let ${(H,\mathcal{B})}$ be an abelian RN hypergroup and $\mu\in{{G{\widehat{(H,\mathcal{B})}}}}$. Then, $\mu\star\mu_{k}\in\widehat{B}$, for any $\mu_{k}\in\widehat{B}$.
@@
Since $\mu\in{{G{\widehat{(H,\mathcal{B})}}}}$ and ${(H,\mathcal{B})}$ is RN, by Lemma 2.28, $\mu(x_{i})=\omega_{i}d_{i}$ for some root of unity $\omega_{i}$ and for all $i\in\mathcal{I}$. By Lemma 2.35, $\mu(x_{l})=\omega_{i}\omega_{j}d_{l}$ for any constituent $x_{l}$ of $x_{i}x_{j}$, therefore
@@
By duality, Lemma 2.36 and Proposition 2.37 reformulate as follows:
@@
If ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN, then $\overline{G{(H,\mathcal{B})}}$ is a group. Moreover, the left multiplication $L_{{\frac{x_{i}}{d_{i}}}}$, for any $x_{i}\in{G{(H,\mathcal{B})}}$, permutes the normalized basis $\overline{\mathcal{B}}$.
Lemmas 2.32, 2.33, 2.34, 2.36 and Propositions 2.37, 2.38 imply the following:
@@
    $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN.
Then, the following statements hold:
@@
    The left multiplication operator $L_{\frac{x_{i}}{d_{i}}}$, with $x_{i}\in G{(H,\mathcal{B})}$, permutes the normalized basis $\overline{\mathcal{B}}$.
@@
    The left multiplication operator $L_{\mu_{i}}$, with $\mu_{i}\in G{\widehat{(H,\mathcal{B})}}$, permutes the basis $^{{\widehat{\mathcal{B}}}}$.
2.6. Some results on $P$ and $^{\widehat{P}}$.
They denote $\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}$ and $\prod_{j\in\mathcal{I}}\mu_{j}$, respectively.
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and let $x_{i}\in G{(H,\mathcal{B})}$ be a grouplike element. Then
(2.28) $(\prod_{j\in\mathcal{I}}\mu_{j})({\frac{x_{i}}{d_{i}}})=\pm 1.$
@@
By Equation (2.7), defining the multiplication in the dual,
$(\prod_{j\in\mathcal{I}}\mu_{j})({\frac{x_{i}}{d_{i}}})=\prod_{j\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}}).$
By Equation (2.8), defining the involution in the dual, and then Lemma 2.27,
$\mu_{j}({\frac{x_{i}}{d_{i}}})\mu_{j^{\#}}({\frac{x_{i}}{d_{i}}})=|\mu_{j}({\frac{x_{i}}{d_{i}}})|^{2}=1.$
We group together the factors $\mu_{j}$ and $\mu_{j\#}$ in the product above. Now, if $\mu_{j}=\mu_{j^{\#}}$, then $\mu_{j}({\frac{x_{i}}{d_{i}}})\in\mathbb{R}$, and therefore, $\mu_{j}({\frac{x_{i}}{d_{i}}})=\pm 1$ by Lemma 2.27. ∎
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and let $x_{i}\in{G{(H,\mathcal{B})}}$ be a grouplike element. Then $(\prod_{j\in\mathcal{I}}\mu_{j})({\frac{x_{i}}{d_{i}}})$ will be denoted ${\mathrm{sgn}}(x_{i})$.
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and let $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ be a grouplike element. Then
(2.29) $\mu_{j}(\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}})=\pm 1.$
By duality, we can also define:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup, and let $\mu_{j}\in G{\widehat{(H,\mathcal{B})}}$ be a grouplike element. Then $\mu_{j}(\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}})$ will be denoted ${\mathrm{sgn}}(\mu_{j})$.
Proposition 2.44.
Following the assumptions of Corollary 2.39, let $x_{i}\in{G{(H,\mathcal{B})}}$ and $\mu_{j}\in G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ be grouplike elements. Then
${\mathrm{sgn}}(x_{i})=\det(L_{\frac{x_{i}}{d_{i}}})\text{ and }{\mathrm{sgn}}(\mu_{j})=\det(L_{\mu_{j}})$
are the signature of the permutations of $\overline{\mathcal{B}}$ by $L_{{\frac{x_{i}}{d_{i}}}}$, and of $\widehat{\mathcal{B}}$ by $L_{\mu_{j}}$, respectively.
Proof.
By Corollary 2.39, the left multiplication operator $L_{{\frac{x_{i}}{d_{i}}}}$ permutes $\overline{\mathcal{B}}$, but the signature of this permutation is the determinant $\det(L_{\frac{x_{i}}{d_{i}}})$. Now ${\frac{x_{i}}{d_{i}}}=\sum_{j\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})F_{j}$, so $\det(L_{\frac{x_{i}}{d_{i}}})=\prod_{j\in\mathcal{I}}\mu_{j}({\frac{x_{i}}{d_{i}}})$. Idem for ${\mathrm{sgn}}(\mu_{j})$. ∎
@@
A hypergroup ${(H,\mathcal{B})}$ is called perfect if it has no nontrivial grouplike elements.